40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 31 23:12:17 2018, gautam, Update, LSC, DRMI locking attempt REFL55_whtCheck.pdf
    Reply  Fri Jun 1 01:22:50 2018, gautam, Update, LSC, DRMI locking restored DRMI_20180531.png
       Reply  Fri Jun 8 01:58:56 2018, gautam, Update, LSC, DRMI locking attempt again 
          Reply  Tue Jun 12 03:22:25 2018, gautam, Update, LSC, AUX laser shuttered 
             Reply  Wed Jun 13 01:02:40 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking 
                Reply  Wed Jun 13 11:17:40 2018, gautam, Update, LSC, PRCL loop shape anomaly PRCL_12Jun2018_WeirdShape.pdf
                   Reply  Thu Jun 14 00:40:42 2018, gautam, Update, LSC, PRCL loop shape anomaly 
                Reply  Thu Jun 14 18:09:24 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking DRMI1f_June14.pdf
                   Reply  Fri Jun 15 00:53:21 2018, gautam, Update, LSC, Calibrated MICH spectrum C1NB_disp_40m_MICH_NB_2018-06-14.pdf
Message ID: 13908     Entry time: Fri Jun 1 01:22:50 2018     In reply to: 13907     Reply to this: 13933
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI locking restored 

As it happened last time, the problem apparently fixed itself - somehow the act of me disconnecting the cables and reconnecting them seems to solve the problem, need to think about this.

Anyway, DRMI was locked a few times tonightyes. I got in a good long stretch where I ran some sensing lines and collected some data, analysis tomorrow. I am going to center the vertex oplevs as an alignment reference for now. A major source of lockloss seems to be angular instability - see for example this video grab of POP:

Could be due to noise injection from the noisy PRM Oplev HeNe, or just TT mirror angular motion (I couldn't get the PRC angular FF going tonight).

Attachment 1: DRMI_20180531.png  32 kB  Uploaded Fri Jun 1 02:29:14 2018  | Show | Show all
ELOG V3.1.3-